OpenStudy (anonymous):

-Can you find the vertex of y = x2 + 3x - 10? I need help .

OpenStudy (anonymous):

somewhere near (-1,-12)

OpenStudy (cwrw238):

divide coefficient of x by 2 this gives 3/2 = 1.5 y = (x + 1.5)^2 - 1.5^2 - 10 y = (x - 1.5)^2 - 12.25 this is the vertex form of equation

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you could just say -b/2a is your x coordinate then plug whatever you get for your x coordinate back into the formula, and that will be your y coordinate (x,y) is the answer.

OpenStudy (cwrw238):

the x coordinate of vertex is in the parentheses and the last number is the vertex or you can do as funinabox said